# Algorithms and Complexity of Hedge Cluster Deletion Problems

### v1

@@ 3 Hardness result and complexity dichotomy
For $\delta\geq 0$, we denote by $\delta P_{3}$ the graph obtained from the union of $\delta$ vertex-disjoint $P_{3}$.
@@
Hedge Cluster Deletion is NP-complete on hedge graphs with underlying graph isomorphic to $\delta P_{3}$, where $n=3\delta$.
@@
  • For each edge $\{x,y\}\in E(G)$, we add three vertices $e_{xy},e_{x}\text{ and }e_{y}$ to $V(H)$. Moreover, we add the edges $\{e_{xy},e_{x}\}$ and $\{e_{xy},e_{y}\}$.
@@
We claim that $G$ has a vertex cover $S$ with $|S|\leq k$ if and only if there are $U$ hedges in $H$ with $|U|\leq k$ such that $H\setminus U$ is a cluster hedge-subgraph of $H$.
@@
Now we are ready to extend our result on a wide range of graph classes. We note that Lemma 3.1 alone is not enough to be applicable on larger underlying graphs (i.e., graphs containing union of $P_{3}$’s), since the rest of the edges need to be included in some of the hedges. The following reduction considers underlying graphs that contain union of $P_{3}$’s without determining the exact structure of the rest of the edges. Let $\delta P_{3}$ be a maximum subgraph of a hedge graph $H$, meaning that the underlying graph of $H$ contains a subgraph (not necessarily induced subgraph) isomorphic to $\delta P_{3}$ with $\delta$ as large as possible. If $\delta$ is unbounded with respect to the size of $H$, we say that $H$ contains a large subgraph $\delta P_{3}$. Otherwise, if $\delta=O(1)$, we say that $H$ contains a small subgraph $\delta P_{3}$.
Proposition 3.2.
Hedge Cluster Deletion is NP-complete on hedge graphs that contain a large subgraph $\delta P_{3}$.
@@
We give a polynomial reduction from the Hedge Cluster Deletion problem on hedge graphs given in Lemma 3.1. Let $H$ be a hedge graph that contains a large subgraph $\delta P_{3}$ and let $H^{\prime}$ be a hedge graph with its underlying graph isomorphic to $\delta P_{3}$. We add all necessary vertices and edges that appear in $H-H^{\prime}$. Moreover, we add three more vertices $x,y,z$ with edges $\{x,y\},\{y,z\}$. For the hedges of $H$, we keep the $\ell$ hedges of $H^{\prime}$ and add an extra hedge $E_{\ell+1}$ that contains all edges of $(E(H)\setminus E(H^{\prime}))\cup\{x,y\}\cup\{y,z\}$. Now we claim that $(H^{\prime},k)$ is a yes-instance if and only if $(H,k+1)$ is a yes-instance for Hedge Cluster Deletion.
@@
If $U^{\prime}$ is a solution for Hedge Cluster Deletion on $H^{\prime}$ then $H\setminus(U^{\prime}\cup\{E_{\ell+1}\})$ is a cluster hedge-subgraph because all vertices of $H-H^{\prime}$ are trivial clusters and $H^{\prime}$ is a cluster hedge-subgraph. For the opposite direction, let $U$ be a solution for Hedge Cluster Deletion on $H$. By 2.4, we have $E_{\ell+1}\subseteq U$ due to the internal $P_{3}$ induced by $\{x,y,z\}$. Thus $U\setminus\{E_{\ell+1}\}$ contains $k$ hedges. Since $H^{\prime}$ is a hedge-subgraph of $H$, we conclude that $H^{\prime}\setminus(U\setminus\{E_{\ell+1}\})$ is a cluster hedge-subgraph. ∎
@@
Typical examples of graph classes that fall in the large subgraph characterization are long paths, and almost cliques (graphs isomorphic to $K_{n}$ without a single edge). Both examples reveal an interesting complexity difference with the classical variation of the problem, because on such graphs Cluster Deletion is known to be solved in polynomial time, whereas by Proposition 3.2 Hedge Cluster Deletion remains NP-complete.
We next conclude the section with a complexity dichotomy result with respect to the $\delta P_{3}$ subgraph characterization. To mention few classes of graphs with small subgraphs isomorphic to $\delta P_{3}$, we note complete bipartite graphs with bounded one side or trees with bounded number of vertex-disjoint $P_{3}$’s.
@@
Hedge Cluster Deletion is polynomial-time solvable on hedge graphs that contain $\delta P_{3}$ for some constant $\delta$. Otherwise, it is NP-complete on hedge graphs that contain a large subgraph $\delta P_{3}$.
@@
The NP-completeness follows by Proposition 3.2. Here we show the claimed polynomial-time algorithm whenever $\delta$ is a fixed small number. First we show how to compute such a subgraph $\delta P_{3}$. We enumerate all subgraphs of $H$ on three vertices in $O(n^{3})$ time. From all three-vertex subgraphs with at least two edges we choose $\delta$ to check if they are vertex-disjoint. In total, we need $O(n^{3\delta})$ time to compute the largest $\delta$ and the corresponding subgraph of $H$.
We now describe how to compute a solution for Hedge Cluster Deletion. In fact, we will show that all possible cluster subgraphs of $H$ can be enumerated in polynomial time. This leads to a polynomial-time algorithm by 2.3. Let $H_{\delta}$ be the subgraph of $H$ that is isomorphic to $\delta P_{3}$. The $3\delta$ vertices of $H_{\delta}$ are denoted by $V_{\delta}$. We partition the vertices of $V(H)\setminus V_{\delta}$ into two sets: $M$ vertices that form a maximum matching in $H-H_{\delta}$ and $Q$ vertices that do not belong to $M$. We claim that there are no other edges in $H[M\cup Q]$ except the ones belonging to the matching. To see this, observe the following:
@@
4 (In)approximability of Hedge Cluster Deletion
@@
We will apply $A$-reductions from the Min Horn Deletion-complete problem MinOnes()($\mathcal{F}$): given a set of Boolean variables and a family of characteristic constraints $\mathcal{F}$ over the variables, find a variable assignment that satisfy all of the constraints while minimizing the number of variables set to one (i.e., variables assigned to true). We refer to characteristic constraints, to mean that every constraint $f$ is constructed from $\mathcal{F}$ by applying certain rules on some tuple of the variables. To be more precise, we define the necessary properties of Boolean constraints.
@@
  • If none of the variables are set to $1$, no hedges are deleted from the corresponding $K_{3}$ in $H$.
@@
This completes the construction of the instance $\phi$ for MinOnes(′′)($\mathcal{F}^{\prime\prime}$). It is not difficult to see that all steps can be performed in polynomial time.
@@
For the opposite direction, assume that the assignment $\{x_{i}=1|i\in I\}$ satisfies $\phi$. We show that every connected component of $H$ is a cluster graph in $H\setminus U$. Constraints $g_{2}$ and $f_{2}$ guarantee that each connected component of $H$ isomorphic to $P_{3}$ in $H$ does not contain at least one of its two edges in $H\setminus U$. For every connected component of $H$ isomorphic to $K_{3}$, we apply constraints $f_{1}$ or $g_{1}$ depending on the number of hedges that covers the $K_{3}$. We note that if the $K_{3}$ is covered by exactly one hedge $X$ then regardless of whether $X\in U$ or not, the three vertices of the connected component induce a cluster graph in $H\setminus U$. If the $K_{3}$ is covered by two hedges $X_{1},X_{2}$ with $X_{1}$ the hedge that covers the unique edge of $K_{3}$, then the constraint $g_{2}$ forces that $X_{1}\in U$ and $X_{2}\notin U$ never occurs. Otherwise, the $K_{3}$ is covered by three hedges and the constraint $f_{1}$ is responsible not to remove exactly one of the two edges. Therefore, $H\setminus U$ is a cluster hedge-subgraph of $H$, as claimed. ∎
@@
4.1 Towards a constant-factor approximation

### v2

@@ 3 Hardness result implied by the underlying structure
For $\delta\geq 0$, we denote by $\delta P_{3}$ the graph obtained from the union of $\delta$ vertex-disjoint $P_{3}$’s.
@@
Hedge Cluster Deletion is NP-complete on hedge graphs with underlying graph isomorphic to $\delta P_{3}$.
@@
  • For each edge $e=\{x,y\}\in E(G)$, we add three vertices $e_{xy},e_{x}\text{ and }e_{y}$ to $V(H)$. Moreover, we add the edges $\{e_{xy},e_{x}\}$ and $\{e_{xy},e_{y}\}$.
@@
We claim that $G$ has a vertex cover $S$ with $|S|\leq k$ if and only if there is a set of $U$ hedges in $H$ with $|U|\leq k$ such that $H\setminus U$ is a cluster hedge-subgraph of $H$.
@@
Now we are ready to extend our result on a wide range of graph classes. We note that Lemma 3.1 alone is not enough to be applicable on larger underlying graphs (i.e., graphs containing union of $P_{3}$’s), since the rest of the edges need to be included in some of the hedges. For instance, connected graphs are not applicable by Lemma 3.1, though they contain a large enough union of $P_{3}$’s.
Observe that for an underlying graph $G$, there are many hedge graphs (exponential in $m$) for which their underlying graph is isomorphic to $G$. We collect all such hedge graphs, as follows. Given a graph $G$, we denote by $\mathcal{H}(G)$ the class of hedge graphs that contains all hedge graphs for which their underlying graph is isomorphic to $G$.
Given two graphs $F$ and $G$, an $F$-packing in $G$ is a collection of pairwise vertex-disjoint copies of $F$ in $G$. An $F$-packing in $G$ is called perfect if it covers all the vertices of $G$. We write $F_{3}$ to denote the graph on three vertices that is connected, i.e., $F_{3}\in\{P_{3},K_{3}\}$. It should be noted that the decision problem of whether a graph $G$ has a perfect $F_{3}$-packing is NP-complete [KirkpatrickH83]. However, there are simple classes of graphs in which a perfect $F_{3}$-packing can be computed efficiently, such as trees or complete graphs.
Now we are ready to state our claimed NP-completeness result. We take advantage of a perfect $F_{3}$-packing of the underlying graph and spread the hedges in a particular way that exhibits the difficulty shown in Lemma 3.1. We stress that we do not claim NP-completeness on any non-cluster graph, but insist on non-cluster graphs for which a perfect $F_{3}$-packing can be computed in polynomial-time.
Theorem 3.2.
For any non-cluster graph $G$ with a perfect $F_{3}$-packing, Hedge Cluster Deletion is NP-complete on the class of $\mathcal{H}(G)$ hedge graphs.
@@
We give a polynomial reduction from the Hedge Cluster Deletion problem on hedge graphs proved to be NP-complete in Lemma 3.1. Let $H^{\prime}$ be a hedge graph with its underlying graph isomorphic to $\delta^{\prime}P_{3}$ having $\ell^{\prime}$ hedges and $n^{\prime}$ vertices. As we claim the construction for any non-cluster underlying graph, we will use $H^{\prime}$ only to define the hedges of the given underlying graph $G$. Let $C_{1},\ldots,C_{\delta}$ be the collection of the three-vertex connected subgraphs in the given perfect $F_{3}$-packing of $G$, where $n=3\delta$. Since $G$ is non-cluster, there is a $P_{3}=(x,y,z)$ in $G$. We denote by $X$ be the set of the connected subgraphs of the collection having at least one vertex from the $P_{3}$, that is, $X=\{C_{i}\mid C_{i}\cap\{x,y,z\}\neq\emptyset\}$. Note that $|V(X)|\in\{3,6,9\}$. Now we describe the hedge graph $H$ such that $\delta=\delta^{\prime}+|X|$ and $n=n^{\prime}+|V(X)|$. Let $C_{j}\in\{C_{1},\ldots,C_{\delta}\}\setminus X$. Observe that $C_{j}$ contains two or three edges. We arbitrarily map two edges of $C_{j}$ into two edges of a connected component of $H^{\prime}$ (recall that $H^{\prime}=\delta^{\prime}P_{3}$). According to the mapping for each $C_{j}$, we construct $\ell^{\prime}$ hedges $E_{1},\ldots,E_{\ell^{\prime}}$ in $H$ that respect the corresponding hedges of $H^{\prime}$. For any other edge of $H$ that it is not mapped, we include it into a single hedge $E_{\ell^{\prime}+1}$. In particular, notice that the two edges of the $P_{3}=(x,y,z)$ belong to $E_{\ell^{\prime}+1}$. Thus $H$ contains $\ell^{\prime}+1$ hedges. Now we claim that $(H^{\prime},k)$ is a yes-instance if and only if $(H,k+1)$ is a yes-instance for Hedge Cluster Deletion.
@@
If $U^{\prime}$ is a solution for Hedge Cluster Deletion on $H^{\prime}$ then $H\setminus(U^{\prime}\cup\{E_{\ell^{\prime}+1}\})$ is a cluster hedge-subgraph because all vertices of $V(X)=H-H^{\prime}$ are trivial clusters and $H^{\prime}$ is a cluster hedge-subgraph. For the opposite direction, let $U$ be a solution for Hedge Cluster Deletion on $H$. By 2.4, we have $E_{\ell^{\prime}+1}\subseteq U$ due to the internal $P_{3}$ induced by $\{x,y,z\}$. Thus $U\setminus\{E_{\ell^{\prime}+1}\}$ contains $k$ hedges. Since $H^{\prime}$ is a hedge-subgraph of $H$, we conclude that $H^{\prime}\setminus(U\setminus\{E_{\ell^{\prime}+1}\})$ is a cluster hedge-subgraph. ∎
@@
As an application of Theorem 3.2, we require hedge graphs for which their underlying graph admits a polynomial-time algorithm to compute a perfect $F_{3}$-packing. Typical examples of graph classes that fall in the previous characterization are long paths and almost cliques (graphs isomorphic to $K_{n}$ without a single edge). Both examples reveal an interesting complexity difference with the classical variation of the problem, because on such graphs Cluster Deletion is known to be solved in polynomial time, whereas by Theorem 3.2 Hedge Cluster Deletion remains NP-complete.
We next conclude the section with a polynomial-time algorithm with respect to the $F_{3}$-packing characterization. The size of an $F_{3}$-packing refers to the number of vertex-disjoint subgraphs in the collection. For a graph $G$, we denote by $\delta_{\max}$ the size of the largest $F_{3}$-packing. If $G$ admits a perfect $F_{3}$-packing then $\delta_{\max}=n/3$, whereas if $G$ is a tree of height one then $\delta_{\max}=1$. For a hedge graph $H$, we also write $\delta_{\max}$ to refer to the actual number of its underlying graph. We show that if $\delta_{\max}$ is constant then Hedge Cluster Deletion admits a polynomial-time solution. In particular, our claimed algorithm has running time $n^{O(\delta_{\max})}$. To mention few classes of graphs with small $\delta_{\max}$, we note complete bipartite graphs with bounded one side or trees with bounded number of vertex-disjoint $P_{3}$’s.
@@
Hedge Cluster Deletion is polynomial-time solvable on hedge graphs for which $\delta_{\max}=O(1)$.
@@
Let $H$ be a hedge graph and let $\delta=\delta_{\max}$. We show the claimed polynomial-time algorithm whenever $\delta$ is a fixed small number. First we show how to compute such a collection of vertex-disjoint connected subgraphs on $3$ vertices of $H$ of size $\delta$. We enumerate all subgraphs of $H$ on three vertices in $O(n^{3})$ time. From all three-vertex subgraphs with at least two edges, we choose $\delta$ to check if they are vertex-disjoint. In total, we need $O(n^{3\delta})$ time to compute the largest $\delta$ and the corresponding $F_{3}$-packing of $H$.
We now describe how to compute a solution for Hedge Cluster Deletion. In fact, we will show that all possible cluster subgraphs of $H$ can be enumerated in polynomial time. This leads to a polynomial-time algorithm by 2.3. Let $H_{\delta}$ be the subgraph of $H$ that is isomorphic to $\delta P_{3}$. By the $F_{3}$-packing in $H$, such a subgraph always exist. The $3\delta$ vertices of $H_{\delta}$ are denoted by $V_{\delta}$. We partition the vertices of $V(H)\setminus V_{\delta}$ into two sets: $M$ the vertices that form a maximum matching in $H-H_{\delta}$ and $Q$ the vertices that do not belong to $M$. We claim that there are no other edges in $H[M\cup Q]$ except the ones belonging to the matching. To see this, observe the following:
@@
4 (In)approximability and Incompressibility
Here we show that Hedge Cluster Deletion is Min Horn Deletion-complete and has no polynomial kernel (or polynomial compression), unless NP $\subseteq$ coNP/poly. When we consider the approximability, we naturally refer to the optimization version of the problem, whereas for the kernelization we refer to the decision version parameterized by the size of the solution (i.e., the number of hedge removals). In both directions we will apply boolean constraint satisfaction problems. Before introducing some relevant notation, let us first describe where the difficulty occurs when approaching the problem in both situations.
It is not difficult to see that Hedge Cluster Deletion is closely related to Vertex Cover (just consider only the $P_{3}$’s in Lemma 3.1). However, it appears that one needs to solve a vertex cover in a “dynamic manner”: whenever a cover solution is found (for $P_{3}$-freeness), its removal results in new $P_{3}$’s that are not nicely structured in the original graph (except the fact that they appear in $K_{3}$’s). In some sense, the obstacles can propagate in the graph under each hedge removal. This chain of implications might result in a solution that is far from optimal and not within a reasonable guarantee. Thus, the main obstacle is that of introducing new induced copies of $P_{3}$ when we delete hedges, which causes a propagation of hedge deletions. This propagation phenomenon also arises in CSP problems when assigning true to Boolean variables in certain formulas that we describe next.
We employ standard notation related to constraint satisfiability problems. A CNF formula $\phi$ is a conjuction of clauses where each clause is a disjunction of literals. A satisfying assignment for $\phi$ is an assignment on its Boolean variables such that each clause in $\phi$ is satisfied. For each Boolean variable, we interchange the value of one to represent true and zero to represent false. The degree of a Boolean variable is the number of its occurrences in $\phi$ and $\phi$ is called 3-regular if all of its variables have degree 3. Moreover, the weight of a truth assignment for $\phi$ is the number of ones in the assignment.
For each clause in $\phi$, we will apply a Boolean constraint or function $f$ as follows. As we deal with 3CNF, that is, each clause is the disjunction of at most three literals, for $f(x,y,z)$ the conjuctive formula $\phi$ is of the form $f(x_{1},y_{1},z_{1})\wedge\cdots\wedge f(x_{m},y_{m},z_{m})$, where each $f(x_{i},y_{i},z_{i})$ is a clause of $\phi$.
4.1 (In)approximability of Hedge Cluster Deletion
@@
Though $A$-reductions preserve approximability problems up to a constant factor (or higher), we will actually apply $A$-reductions as natural polynomial-time reductions. In fact we show that there is a polynomial-time reduction from a Min Horn Deletion-complete problem to Hedge Cluster Deletion (and vice versa), that preserves the same size of the solutions. Our starting point is the Min Horn Deletion-complete problem MinOnes()($\mathcal{F}$): given a set of boolean variables and a family of characteristic constraints $\mathcal{F}$ over the variables, find a variable assignment that satisfy all of the constraints while minimizing the number of variables set to one (i.e., variables assigned to true). We refer to characteristic constraints, to mean that every constraint $f$ is constructed from $\mathcal{F}$ by applying certain rules on some tuple of the variables. To be more precise, we define the necessary properties of boolean constraints.
@@
  • If none of the variables are set to $1$, no hedge is deleted from the corresponding $K_{3}$ in $H$.
@@
This completes the construction of the instance $\phi$ for MinOnes(′′)($\mathcal{F}^{\prime\prime}$) (see also Figure 2). It is not difficult to see that all steps can be performed in polynomial time.
@@
For the opposite direction, assume that the assignment $\{x_{i}=1\mid i\in I\}$ satisfies $\phi$. We show that every connected component of $H$ is a cluster graph in $H\setminus U$. Constraints $g_{2}$ and $f_{2}$ guarantee that each connected component of $H$ isomorphic to $P_{3}$ in $H$ does not contain at least one of its two edges in $H\setminus U$. For every connected component of $H$ isomorphic to $K_{3}$, we apply constraints $f_{1}$ or $g_{1}$ depending on the number of hedges that covers the $K_{3}$. We note that if the $K_{3}$ is covered by exactly one hedge $X$ then regardless of whether $X\in U$ or not, the three vertices of the connected component induce a cluster graph in $H\setminus U$. If the $K_{3}$ is covered by two hedges $X_{1},X_{2}$ with $X_{1}$ the hedge that covers the unique edge of $K_{3}$, then the constraint $g_{2}$ forces that $X_{1}\in U$ and $X_{2}\notin U$ never occurs. Otherwise, the $K_{3}$ is covered by three hedges and the constraint $f_{1}$ is responsible not to remove exactly one of the two edges. Therefore, $H\setminus U$ is a cluster hedge-subgraph of $H$, as claimed. ∎
@@
4.2 Incompressibility of Hedge Cluster Deletion
Recall that with a simple brute-force algorithm Hedge Cluster Deletion can be solved in $2^{k}\cdot n^{O(1)}$ time. Here we show that Hedge Cluster Deletion does not have a polynomial kernel parameterized by $k$ unless the polynomial hierarchy collapses.
A ternary boolean function $f(x,y,z)$, where $x,y$ and $z$ are either boolean variables or constants 0 or 1, is propagational if it satisfies
  • $f(1,0,0)=0$ and
  • $f(0,0,0)=f(1,0,1)=f(1,1,0)=f(1,1,1)=1$.
It is not difficult to see that the following Boolean function is propagational [CaiC15, KratschW13]:
$\texttt{Not-1-in-3}(x,y,z)=f_{1}(x,y,z)=(\lnot x\lor y\lor z)\land(x\lor\lnot y\lor z)\land(x\lor y\lor\lnot z).$
Note that $f_{1}$ is symmetric ternary function and the order of the three arguments in $f_{1}$ is not important. Within the list of arguments, we write first the constants (i.e., $f_{1}(0,x,y)=f_{1}(x,0,y)$, $f_{1}(0,1,x)=f_{1}(1,0,x)$). Also notice that $f_{1}\in\mathcal{F^{\prime}}\cap\mathcal{F^{\prime\prime}}$ in our previous considered MinOnes problem which was used to encode every $K_{3}$ that contains three hedges.
We will reduce from the Propagational-$f$ Satisfiability that is defined as follows: given a conjunctive formula $\phi$ of a propagational ternary function $f$ with distinct variables inside each clause of $\phi$ and an integer parameter $k\geq 0$, the task is to decide whether there exists a satisfying truth assignment for $\phi$ of weight at most $k$. Cai and Cai [CaiC15] introduced the problem and proved that it has no polynomial kernel under plausible hierarchy assumptions.
Theorem 4.10 ([CaiC15]).
For any propagational ternary boolean function $f$, Propagational-$f$ Satisfiability on 3-regular conjunctive formulas admits no polynomial compression, hence no polynomial kernel, unless NP $\subseteq$ coNP/poly.
In what follows, we set $f_{1}$ as a propagational function and consider the Propagational-$f_{1}$ Satisfiability problem. For technical reasons, it is more convenient to deal with $\phi$ that does not contain a function of the form $f_{1}(0,0,x)$. Next we show how this assumption is achieved.
Lemma 4.11.
Let $(\phi,k)$ be an instance of Propagational-$f_{1}$ Satisfiability. In polynomial time we can compute a Boolean formula $\phi^{\prime}$ from $\phi$ such that $f_{1}(0,0,x)\notin\phi^{\prime}$ and $(\phi,k)$ is a yes-instance if and only if $(\phi^{\prime},k)$ is a yes-instance.
Proof.
We construct $\phi^{\prime}$ from $\phi$ by removing certain clauses. For each Boolean variable $x$, we check whether $f_{1}(0,0,x)\in\phi$. If this is not the case for every $x$, then $\phi^{\prime}=\phi$ and we conclude with the instance $(\phi^{\prime},k)$. Assume that $f_{1}(0,0,x)\in\phi$. By the definition of $f_{1}$, observe that $\phi$ is satisfied only if $x=0$. We construct $\phi^{\prime}$ from $\phi$ by replacing every occurrence of $x$ with the constant 0 in $\phi$. Since we have not assigned one to any variable during this construction, it is not difficult to see that $\phi$ is satisfied by at most $k$ ones if and only if $\phi^{\prime}$ is satisfied by at most $k$ ones. Notice also that clauses in $\phi^{\prime}$ that contain only constants can be safely removed, because if we encounter $f_{1}(0,0,1)\in\phi^{\prime}$ then $(\phi,k)$ is a no-instance; otherwise we just ignore such clauses (as they are always true). By applying such a transformation for every variable, we conclude with a valid $\phi^{\prime}$ such that $f_{1}(0,0,x)\notin\phi^{\prime}$. ∎
We next establish the claimed incompressibility of Hedge Cluster Deletion mainly by a ppt-reduction from Propagational-$f_{1}$ Satisfiability. The proof is similar to the reduction given in Lemma 4.4.
[…]
Figure 2: Illustrating CNF formulas for MinOnes(′′)($\mathcal{F}^{\prime\prime}$) and Propagational-$f_{1}$ Satisfiability, alongside the corresponding hedge graph given in Lemma 4.6 and Theorem 4.12, respectively. Note that the $K_{3}$ spanned solely by hedge $x_{6}$} does not correspond to any constraint in the MinOnes(′′)($\mathcal{F}^{\prime\prime}$) formulation.
Theorem 4.12.
Hedge Cluster Deletion parameterized by the solution size admits no polynomial compression, hence no polynomial kernel, unless NP $\subseteq$ coNP/poly.
Proof.
We give a ppt-reduction from Propagational-$f_{1}$ Satisfiability to the parameterized problem of Hedge Cluster Deletion. Let $(\phi,k)$ be an instance of Propagational-$f_{1}$ Satisfiability. By Lemma 4.11, we assume that $f_{1}(0,0,x)\notin\phi$. As $f_{1}$ is symmetric ternary and propagational function, the following forms can only occur in $\phi$: $f_{1}(0,1,x)$, $f_{1}(1,1,x)$, $f_{1}(0,x,y)$, $f_{1}(1,x,y)$, $f_{1}(x,y,z)$, where $x,y,z$ are Boolean variables. We construct a hedge graph $H$ from $\phi$ in which each hedge $E_{x}$ of $H$ corresponds to exactly one Boolean variable $x$ of $\phi$. For each clause in $\phi$, we construct some connected components in $H$ where every connected component has exactly three vertices in $H$ (so that every component is a $P_{3}$ or a $K_{3}$). Depending on the five formulations of $f_{1}$, we construct vertex-disjoint connected components as follows:
  • $f_{1}(0,1,x)$: add a $P_{3}$ that is covered by exactly one hedge $E_{x}$.
  • $f_{1}(1,1,x)$: add a $K_{3}$ that is covered by exactly one hedge $E_{x}$.
  • $f_{1}(0,x,y)$: add two $K_{3}$’s where the edges of the two components are covered by exactly two hedges $E_{x}$ and $E_{y}$, such that the two edges of the one component belong to $E_{x}$ and the two edges of the other component belong to $E_{y}$.
  • $f_{1}(1,x,y)$: add a $P_{3}$ that is covered by exactly two hedges $E_{x},E_{y}$.
  • $f_{1}(x,y,z)$: add a $K_{3}$ that is covered by exactly three hedges $E_{x},E_{y},E_{z}$.
This concludes the construction of $H$ (see also Figure 2). It is not difficult to see that the construction can be performed in polynomial time. Now we claim that $(\phi,k)$ is a yes-instance of Propagational-$f_{1}$ Satisfiability if and only if $(H,k)$ is a yes-instance of Hedge Cluster Deletion.
Let $X^{\prime}$ be a subset of the Boolean variables such that $\{x=1\mid x\in X^{\prime}\}$ satisfies $\phi$. We show that $H\setminus E(X^{\prime})$ is a cluster hedge-subgraph where $E(X^{\prime})=\{E_{x}\mid x\in X^{\prime}\}$. Assume for contradiction that there is $P_{3}=(u,v,w)$ in $H\setminus E(X^{\prime})$. This means that the hedges $E_{x}$ and $E_{y}$ of the edges $\{u,v\}$ and $\{v,w\}$, respectively, do not belong to $E(X^{\prime})$, so that for the corresponding variables we have $x=y=0$ (notice that $E_{x},E_{y}$ are not necessarily different hedges). Moreover, if $\{u,w\}$ is an edge of $H$ then we know that $E_{z}\notin\{E_{x},E_{y}\}$ where $E_{z}$ is the hedge of the edge $\{u,w\}$ and for its corresponding variable we have $z=1$. Depending on whether $E_{x}=E_{y}$ and $\{u,w\}\in E(H)$, we conclude that a particular clause in $\phi$ is not satisfied:
  • $E_{x}=E_{y}$ and $\{u,w\}\notin E(H)$. Then $f_{1}(0,1,x)=0$, because $x=0$;
  • $E_{x}=E_{y}$ and $\{u,w\}\in E(H)$. Then $f_{1}(0,x,z)=0$, because $x=0$ and $z=1$;
  • $E_{x}\neq E_{y}$ and $\{u,w\}\notin E(H)$. Then $f_{1}(1,x,y)=0$, because $x=y=0$;
  • $E_{x}\neq E_{y}$ and $\{u,w\}\in E(H)$. Then $f_{1}(x,y,z)=0$, because $x=y=0$ and $z=1$.
Therefore, in each case we conclude with a clause that is not satisfied, leading to a contradiction, which implies that there is no $P_{3}$ in $H\setminus E(X^{\prime})$.
For the opposite direction, let $H^{\prime}=H\setminus E(X^{\prime})$ be a cluster hedge-subgraph where $E(X^{\prime})$ is a subset of the hedges of $H$. Then we claim that $\{x=1\mid x\in E(X^{\prime})\}$ satisfies $\phi$. We show that in each formulation of $f_{1}$ in $\phi$, $f_{1}(x,y,z)$ is satisfied.
  • $f_{1}(0,1,x)$: the corresponding $P_{3}$ is not present in $H^{\prime}$, so that $x\in E(X^{\prime})$ and $x=1$ which means that $f_{1}(0,1,x)$ is satisfied.
  • $f_{1}(1,1,x)$: there are two cases to consider in $H^{\prime}$. Either the corresponding $K_{3}$ remains the same ($x=0$) or the three vertices belong to singleton clusters ($x=1$). Thus, $f_{1}(1,1,x)$ is satisfied.
  • $f_{1}(0,x,y)$: there are two cases to consider in $H^{\prime}$. Either the two corresponding components are untouched ($x=0$ and $y=0$) or all six vertices belong to singleton clusters ($x=1$ and $y=1$). Thus, $f_{1}(0,x,y)$ is satisfied.
  • $f_{1}(1,x,y)$: the corresponding $P_{3}$ is not present in $H^{\prime}$, so that $x\in E(X^{\prime})$ and $x=1$ or $y\in E(X^{\prime})$ and $y=1$ (or both belong to $E(X^{\prime})$ and $x=y=1$). Thus, $f_{1}(1,x,y)$ is satisfied.
  • $f_{1}(x,y,z)$: either all edges of the corresponding $K_{3}$ are present in $H^{\prime}$ ($x=y=z=0$) or at least two of its edges are removed ($x=y=1$). Thus, $f_{1}(x,y,z)$ is satisfied.
Therefore, all clauses of $\phi$ are satisfied as claimed. ∎
5 Towards a constant-factor approximation